OpenStudy (anonymous):

Angle C is an inscribed angle of circle P. angle C measures (-x+8) and arc AB measures (12x) Find X

OpenStudy (anonymous):



OpenStudy (anonymous):

@only4u

OpenStudy (nurali):

(12x)/2 = -x+8 6x= -x+8 6x+x = 8 7x = 8 x=8/7

OpenStudy (anonymous):

Once again @Nurali, Thanks :)

OpenStudy (nurali):

My Pleasure.